# Privacy Issues in Large Language Models: A Survey

### v3

@@ 2 General Terminology
Acronym
Explanation
AUC
Area Under the Curve
BERT
Bidirectional Encoder Representations from Transformers
BLEU
Bilingual Evaluation Understudy
BPE
Byte-Pair Encoding
CRT
Confidentially Redacted Training
DP
Differential Privacy
DP-SGD
Differentially Private Stochastic Gradient Descent
GPT
Generative Pre-trained Transformer
GPT-n, ChatGPT, …
Variants of above
$k$-NAF
$k$ Near Access Freedom
KNN
$k$-Nearest Neighbors
LM
Language Model
LLM
Large Language Model
LOO
Leave One Out
LSTM
Long Short Term Memory network
MIA
Membership Inference Attack
MLM
Masked Language Model
PHI
Personal Health Information
PII
Personally Identifiable Information
SGD
Stochastic Gradient Descent
SISA
Sharded, Isolated, Sliced, and Aggregated training
TPR, FPR
True Positive Rate, False Positive Rate
@@
Dataset
Description
Citation
AG News Corpus
News articles
Gulli (2004)
BigPatent-G
US Patent documents
Sharma et al. (2019)
BookCorpus
Collection of free novel books
Zhu et al. (2015)
Brown
Modern English linguistic corpus
Francis and Kucera (1979)
CNNDM
Reading comprehension data
Hermann et al. (2015)
Common Crawl
Data scraped from open web
Crawl (2008)
C4
Cleaned version of Common Crawl
Dodge et al. (2021)
mC4
Multi-lingual & cleaned Common Crawl
Xue et al. (2021)
Conf. on Machine Translation
Corpora obtained from this conference work
Bojar et al. (2018)
CustomerSim
Synthetic dialogue of customer interactions
Tkachenko et al. (2016)
MIMIC-III
Patient data from intensive care units
Johnson et al. (2016)
One Billion Word
Preprocessed data from 2011 news
Chelba et al. (2014)
OpenWebText
Web content extracted from Reddit URLs
Gokaslan and Cohen (2019)
Penn TreeBank
Annotated corpus of English language
Marcus et al. (1993)
The Pile
Diverse compilation of $22$ smaller datasets
Gao et al. (2020)
PubMed
Citations/abstracts of biomedical literature
Peng et al. (2019)
RoBERTa
Corpus used to train RoBERTa model
Liu et al. (2019)
Sentiment140
Collection of tweets from X
Go (2009)
Stack Overflow Q&A
Data collected from Stack Overflow
Overflow (2019)
TLDR News
Daily tech newsletter collection
Belvèze (2022)
TREC
Q&A dataset
Voorhees and Tice (2000)
Wikicorpus
Wikipedia in English, Spanish, and Catalan
Reese et al. (2010)
Wikipedia
XML text dump of English Wikipedia
Mahoney (2009)
Wikitext-103
Wikipedia articles
Merity et al. (2016)
Wikitext-2
Wikipedia articles
Merity et al. (2016)
DBPedia
Structured Wikipedia content
Zhang et al. (2016a)
XSum
Summaries of news articles
Narayan et al. (2018)
@@
Act
Acronym
Jurisdiction
Year
General Data Privacy Regulation
GDPR
European Union
2018
California Consumer Privacy Act
CCPA
California
2018
Consumer Privacy Protection Act
CPPA
Canada
In progress
New York Privacy Act
NYPA
New York
In progress
@@ 3.1 Preliminaries
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 3.1.1 Additional Definitions
Entity Memorization: When measuring language model memorization, using counterfactual memorization can be computationally expensive while eidetic memorization requires the adversary to have access to the original training data to prompt the prefixes. Zhou et al. (2023) defined another memorization measure, known as entity memorization.
@@
(Entity Memorization, Zhou et al. (2023)) A training string $s$ contains $m$ entities $M$ and can be uniquely identified by any $n$ entities $N$ of them. $n\in(0,m]$. A prompt $c$ contains the $n$ entities that could identify $s$, and $c$ expects an entity $E\in(M-N)$. Prompted the language model $h$ with $c$. If the expected entity $E$ is a substring in the output $h(c)$, the process is referred to as Entity Memorization.
@@
Figure 1: From Zhou et al. (2023). A comparison of typical verbatim memorization versus entity memorization (Definition 8). Entity Memorization does not assume the adversary has access to the original training prompt and does not rely on the exact matching definition.
@@ 4.1 Preliminaries
[…]
@@
[…]
@@ 4.2.2 Shadow Model Attacks
Shadow Models
k
TPR @ $0.1\%$ FPR
TPR @ $1\%$ FPR
$16$
$\sqrt{|D_{meta}|}$
$1.6\%$
$5.2\%$
$16$
$16$
$2.6\%$
$7.0\%$
$32$
$\sqrt{|D_{meta}|}$
$2.0\%$
$5.5\%$
$32$
$32$
$3.1\%$
$8.1\%$
$64$
$\sqrt{|D_{meta}|}$
$2.2\%$
$6.0\%$
$64$
$64$
$3.4\%$
$8.8\%$
@@ 5.1.1 Differential Privacy
[…]
@@
[…]
@@
[…]
@@ 5.2.4 Federated Learning
model
data
$\sigma$
$S$
$K$
$\tilde{C}$
$\epsilon$
AccT1
$0.000$
$\infty$
$763430$
$100$
$\infty$
$17.62\%$
$0.003$
15
$763430$
$5000$
$4.634$
$17.49\%$
$0.006$
10
$763430$
$1667$
$2.314$
$17.04\%$
$0.012$
15
$763430$
$1250$
$2.038$
$16.33\%$
$0.003$
15
$10^{8}$
$5000$
$1.152$
$17.49\%$
$0.006$
10
$10^{8}$
$1667$
$0.991$
$17.04\%$
$0.012$
15
$10^{8}$
$1250$
$0.987$
$16.33\%$
@@ 7 Machine Unlearning
This creates a difficult situation for model developers. The most naive solution, known as Leave One Out (LOO) retraining, would just delete an individual’s data and retrain the whole model. However, modern LLMs can take several weeks and millions of dollars to retrain from scratch making this solution infeasible in practice. According to Help Net Security, consumer data deletion requests have increased by $74\%$ from 2021 to 2022 Security (2023). If models were required to be completely retrained each time one of these requests came in, they would be impossible to maintain, especially given the size of the regions with “Right to Be Forgotten” laws either in place or proposed. This has lead to a new field of research called machine unlearning. This field looks to delete an individual’s data and influence from a model using significantly fewer computational resources than retraining it. A successful unlearning algorithm should produce a model that is approximately indistinguishable from the model produced by retraining. In this section, we discuss the general principles of machine unlearning and current developments within the space of language models. For a more comprehensive survey of machine unlearning, see Nguyen et al. (2022).
@@ 7.1 Preliminaries
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 7.3 Unlearning in LLMs
While work on applying machine unlearning to LMs is limited, there are a couple of notable works in the space. Kumar et al. (2022) expands upon the SISA framework for Bourtoule et al. (2020). While the SISA framework showed computational improvements over naive deletion and retraining, it is impractical for application to LLMs since the number of model parameters and data points makes it extremely computationally expensive. Kumar et al. (2022) proposed two variants of SISA that improve it for LLMs. The first is called SISA-FC. Since saving a checkpoint for every slice during the constituent model training is impossible for large text corpora, this framework starts with an LM trained on a generic corpus. The SISA framework is then only used for the task-specific fine-tuning process. While SISA-FC reduces the retraining time and computational storage, model performance generally performs better if the model is fine-tuned for the task the entire time. The second variant, known as SISA-A, addresses this by using the SISA framework for the entire training. To save memory, the SISA-A framework trains the model using the parameter efficient adapter method from Houlsby et al. (2019). Adapter modules are used to replace the model parameters and are significantly smaller, using $1-5\%$ the memory that storing all the model parameters does. This significantly reduces the memory cost of checkpointing during the SISA framework. Kumar et al. (2022) tested these two frameworks on three Glue tasks/corpora Wang et al. (2018): SST-2 Socher et al. (2013), QQP Wang et al. (2018), and MNLI Williams et al. (2018). They trained a BERT model using the different approached and measured the accuracy on these tasks, the time it took to retrain the model after a deletion request, and the amount of memory the final model takes up. SISA-A saw a small dip ($1-6\%$) in performance on the tasks but also saw a 100×100$\times$100 × improvement in retraining time. SISA-FC saw a $30\%$ dip in performance on the tasks compared to SISA-A but also was about 100×100$\times$100 × faster to retrain than SISA-A, making it way faster than the baseline approach. Exact memory numbers were not provided but Kumar et al. (2022) stated that both frameworks saw signficant reductions in memory usage. Since the SISA framework eliminates any influence the requested deleted point may have had, this approach satisfies the exact unlearning definition.
@@
Meanwhile, Jang et al. (2023) studied a gradient based solution for machine unlearning in LMs, named knowledge unlearning. Specifically, given $h_{\theta}$ and a given sequence of tokens that we would like to unlearn, $\textbf{d}=(d_{1},...,d_{T})$, $\mathcal{U}$ solves a gradient ascent problem maximizing the loss function:
@@
Additional Remarks on Unlearning:
Zanella-Béguelin et al. (2020) found that model updates (such as unlearning) can possibly pose a privacy risk if an adversary has access to the versions of the model pre and post update. Through a differential analysis of the model output, a powerful membership inference attack can be performed on the model. For more details on this, see Section 4. Their attack highlights how machine unlearning is not an orthogonal concept to differential privacy, and can in some cases increase risk from a data privacy perspective.

### v4

@@ 2 General Terminology
Acronym Explanation
AUC Area Under the Curve
BERT Bidirectional Encoder Representations from Transformers
BLEU Bilingual Evaluation Understudy
BPE Byte-Pair Encoding
CRT Confidentially Redacted Training
DP Differential Privacy
DP-SGD Differentially Private Stochastic Gradient Descent
GPT Generative Pre-trained Transformer
GPT-n, ChatGPT, … Variants of above
$k$-NAF $k$ Near Access Freedom
KNN $k$-Nearest Neighbors
LM Language Model
LLM Large Language Model
LOO Leave One Out
LSTM Long Short Term Memory network
MIA Membership Inference Attack
MLM Masked Language Model
PHI Personal Health Information
PII Personally Identifiable Information
SGD Stochastic Gradient Descent
SISA Sharded, Isolated, Sliced, and Aggregated training
TPR, FPR True Positive Rate, False Positive Rate
@@
Dataset Description Citation
AG News Corpus News articles Gulli (2004)
BigPatent-G US Patent documents Sharma et al. (2019)
BookCorpus Collection of free novel books Zhu et al. (2015)
Brown Modern English linguistic corpus Francis and Kucera (1979)
CNNDM Reading comprehension data Hermann et al. (2015)
Common Crawl Data scraped from open web Crawl (2008)
C4 Cleaned version of Common Crawl Dodge et al. (2021)
mC4 Multi-lingual & cleaned Common Crawl Xue et al. (2021)
Conf. on Machine Translation Corpora obtained from this conference work Bojar et al. (2018)
CounterFact Factual completion prompts Meng et al. (2022)
CustomerSim Synthetic dialogue of customer interactions Tkachenko et al. (2016)
LIMA Forums and StackExchange data Zhou et al. (2023a)
MIMIC-III Patient data from intensive care units Johnson et al. (2016)
One Billion Word Preprocessed data from 2011 news Chelba et al. (2014)
OpenWebText Web content extracted from Reddit URLs Gokaslan and Cohen (2019)
Penn TreeBank Annotated corpus of English language Marcus et al. (1993)
The Pile Diverse compilation of $22$ smaller datasets Gao et al. (2020)
PubMed Citations/abstracts of biomedical literature Peng et al. (2019)
RoBERTa Corpus used to train RoBERTa model Liu et al. (2019)
Sentiment140 Collection of tweets from X Go (2009)
Stack Overflow Q&A Data collected from Stack Overflow Overflow (2019)
TLDR News Daily tech newsletter collection Belvèze (2022)
TREC Q&A dataset Voorhees and Tice (2000)
Wikicorpus Wikipedia in English, Spanish, and Catalan Reese et al. (2010)
Wikipedia XML text dump of English Wikipedia Mahoney (2009)
Wikitext-103 Wikipedia articles Merity et al. (2016)
Wikitext-2 Wikipedia articles Merity et al. (2016)
DBPedia Structured Wikipedia content Zhang et al. (2016a)
XSum Summaries of news articles Narayan et al. (2018)
zsRE Question-answer pairs from Wikipedia Levy et al. (2017)
@@
Act Acronym Jurisdiction Year
General Data Privacy Regulation GDPR European Union 2018
California Consumer Privacy Act CCPA California 2018
Consumer Privacy Protection Act CPPA Canada In progress
New York Privacy Act NYPA New York In progress
@@ 3.1 Preliminaries
$\ell\textit{perp}_{\theta}(s)$ $=-\log_{2}\text{Pr}(s_{1}\ldots s_{n}|h_{\theta})$
$=\sum_{i=1}^{n}\left(-\log_{2}\text{Pr}(s_{i}|h_{\theta}(s_{1}\ldots s_{i-1}))\right)$
@@
$\text{rank}_{\theta}(s)=\lvert\{r^{\prime}\in\mathcal{R}:\ell\textit{perp}_{\theta}(r)\leq\ell\textit{perp}_{\theta}(s)\}\rvert$
@@
$\text{exposure}_{\theta}(s)$ $=\log_{2}\lvert\mathcal{R}\rvert-\log_{2}\left(\text{rank}_{\theta}(s)\right)$
$=-\log_{2}\text{Pr}_{r\in\mathcal{R}}\big{[}\ell\textit{perp}_{\theta}(r)\leq\ell\textit{perp}_{\theta}(s)\big{]}$
@@
$\approx-\log_{2}\text{Pr}_{r\in\mathcal{S}}\big{[}\ell\textit{perp}_{\theta}(r)\leq\ell\textit{perp}_{\theta}(s)\big{]}$
@@ 3.1.1 Additional Definitions
Entity Memorization: When measuring language model memorization, using counterfactual memorization can be computationally expensive while eidetic memorization requires the adversary to have access to the original training data to prompt the prefixes. Zhou et al. (2023b) defined another memorization measure, known as entity memorization.
@@
(Entity Memorization, Zhou et al. (2023b)) A training string $s$ contains $m$ entities $M$ and can be uniquely identified by any $n$ entities $N$ of them. $n\in(0,m]$. A prompt $c$ contains the $n$ entities that could identify $s$, and $c$ expects an entity $E\in(M-N)$. Prompted the language model $h$ with $c$. If the expected entity $E$ is a substring in the output $h(c)$, the process is referred to as Entity Memorization.
@@
Figure 1: From Zhou et al. (2023b). A comparison of typical verbatim memorization versus entity memorization (Definition 8). Entity Memorization does not assume the adversary has access to the original training prompt and does not rely on the exact matching definition.
@@ 4.1 Preliminaries
If, for example, the test statistic is loss, then this attack simply classifies the low loss points as in and high loss points as out. The choice of $\tau$ determines the trade-off between the precision and recall of the MIA. If $\tau$ is very high, then the MIA will have a high recall, but low precision rate in identifying members of $D_{train}$. Meanwhile if $\tau$ is low, then it will have a high precision rate, but low recall. This also enables statistics common in machine learning such as the AUC. Loss is typically the first choice based on the principle that models tend to overfit on training data. However, Yeom et al. (2018) found that overfitting, while sufficient, is not necessary for a MIA to be successful and other test statistics may instead be used for a successful MIA. This is especially relevant in the context of LLMs which have much larger training datasets and are thus trained using a single-pass SGD. This means they overfit less making loss-based threshold attacks less effective. A more commonly used test statistic is perplexity, which colloquially measures how surprised the model is to see a given sequence. This is very similar to $\ell\textit{perp}_{\theta}$, defined in the Memorization section (Definition 3).
@@
$\textit{perp}_{\theta}=\text{exp}\Big{(}-\frac{1}{n}\sum_{i=1}^{n}\log_{2}\text{Pr}(s_{i}|h_{\theta}(s_{1}\ldots s_{i-1})\Big{)}$
@@ 4.2.2 Shadow Model Attacks
Shadow Models k TPR @ $0.1\%$ FPR TPR @ $1\%$ FPR
$16$ $\sqrt{|D_{meta}|}$ $1.6\%$ $5.2\%$
$16$ $16$ $2.6\%$ $7.0\%$
$32$ $\sqrt{|D_{meta}|}$ $2.0\%$ $5.5\%$
$32$ $32$ $3.1\%$ $8.1\%$
$64$ $\sqrt{|D_{meta}|}$ $2.2\%$ $6.0\%$
$64$ $64$ $3.4\%$ $8.8\%$
@@ 5.1.1 Differential Privacy
$\text{Pr}(\mathcal{A}(D)\in W)\leq e^{\epsilon}\text{Pr}(\mathcal{A}(D^{\prime})\in W)$
@@
$\text{Pr}(\mathcal{A}(D)\in W)\leq e^{\epsilon}\text{Pr}(\mathcal{A}(D^{\prime})\in W)+\delta$
@@
$\text{log}(\frac{\text{Pr}(h(x;A(D))\in Y)}{\text{Pr}(h(x;A(D^{\prime}))\in Y)})\leq\epsilon$
@@ 5.2.4 Federated Learning
model data
$\sigma$ $S$ $K$ $\tilde{C}$ $\epsilon$ AccT1
$0.000$infty $763430$ $100$infty $17.62\%$
$0.003$ 15 $763430$ $5000$ $4.634$ $17.49\%$
$0.006$ 10 $763430$ $1667$ $2.314$ $17.04\%$
$0.012$ 15 $763430$ $1250$ $2.038$ $16.33\%$
$0.003$ 15 $10^{8}$ $5000$ $1.152$ $17.49\%$
$0.006$ 10 $10^{8}$ $1667$ $0.991$ $17.04\%$
$0.012$ 15 $10^{8}$ $1250$ $0.987$ $16.33\%$
@@
Another approach in the family of federated learning frameworks is local differential privacy (LDP) Kasiviswanathan et al. (2010). Under this framework, individuals perturb their data locally before sharing it with a centralized server for use. LDP offers several advantages over a standard DP-SGD fine-tuning setup: i) There is no requirement for the central server to be trusted and ii) the computational overhead is decreased. One way to achieve LDP is to “sanitize” the text by token-wise redaction before using it for fine-tuning. However, Qu et al. (2021) found that doing this at a token-level loses meaningful context resulting in significant loss of utility. Du et al. (2023a) takes a different approach to LDP to significantly improve utility. First, they relax the standard LDP requirements by using the more generalized metric-LDP Alvim et al. (2018), which tunes the noise added to the output based on the similarity of two inputs given a distance metric. Next, sanitizing is done on sentence embeddings rather than token embeddings, which maintains context while having fewer dimensions. Du et al. (2023a) develops two metric-LDP approaches for this by adapting the Purkayastha mechanism Weggenmann and Kerschbaum (2021) and generalized planar Laplace Wu et al. (2017) method. To further reduce the “curse of dimensionality”, Du et al. (2023a) adapts the random projection map of Wu et al. (2017) to NLP pipelines. Finally, Du et al. (2023a) also allows for optional label sanitizing using randomized response at only minor increase to the privacy budget.
Du et al. (2023a) tested this method using a BERT language model on three different tasks: SST-2, IMDb, and QNLI. In all of these, the authorship was the sensitive information. The baseline, non-DP model achieved around $90\%$ accuracy on all three of these tasks. Using metric-LDP without the dimensionality improvements yielded only around $50\%$ accuracy, but the combination of all the methods yielded accuracy with only approximately $1\%$ drops in accuracy. Du et al. (2023a) tested the privacy guarantees across several different privacy attacks. MIA threshold attacks were only slightly better than guessing, embedding inversion attacks had very low recovery rate, and attribute inference attacks were $40\%$ less successful than on the non-DP model.
The same author group formulated another approach that naturally fits in the federated learning framework. Du et al. (2023b) proposed a method called DP-Forward which directly perturbs embedding matrices in the forward pass of LMs rather than adding noise during the back propagation step. The motivation is that while adding noise during back propagation protects well against membership inference, it is ineffective against inference-time attacks such as sensitive attribute inference and embedding inversion. Furthermore, there are large computational costs in the fine-tuning process. In order to add the minimal amount of noise while maintaining privacy guarantees in the forward pass, Du et al. (2023b) devised the analytical matrix Gaussian mechanism (aMGM). This mechanism exploits a necessary and sufficient DP condition from the analytical Gaussian mechanism by Balle and Wang (2018) to draw possibly non-i.i.d. noise from a matrix Gaussian distribution Chanyaswad et al. (2018). Particularly when dealing with high-dimensional data, this significantly reduces the amount of noise used. Similar to Du et al. (2023a), Du et al. (2023b) tested this using a BERT language model on three different tasks: SST-2, IMDb, and QQP, with authorship being the sensitive information. Compared to the baseline DP-SGD solutions of Yu et al. (2022) and Yu et al. (2021), DP-Forward achieved $7.7\%$ better utility on these tasks while using $3$x less time and memory. The utility of DP-Forward was similar to the non-DP model and provided an $88\%$ reduction in embedding inversion success rate and $41\%$ reduction in sensitive attribute inference success. The baseline DP-SGD was unable to protect against these two inference-time attacks.
@@ 7 Machine Unlearning
This creates a difficult situation for model developers. The most naive solution, known as Leave One Out (LOO) retraining, would just delete an individual’s data and retrain the whole model. However, modern LLMs can take several weeks and millions of dollars to retrain from scratch making this solution infeasible in practice. According to Help Net Security, consumer data deletion requests have increased by $74\%$ from 2021 to 2022 Security (2023). If models were required to be completely retrained each time one of these requests came in, they would be impossible to maintain, especially given the size of the regions with “Right to Be Forgotten” laws either in place or proposed. This has lead to a new field of research called machine unlearning. This field looks to delete an individual’s data and influence from a model using significantly fewer computational resources than retraining it. A successful unlearning algorithm should produce a model that is approximately indistinguishable from the model produced by retraining. In this section, we discuss the general principles of machine unlearning and current developments within the space of language models.
For a more comprehensive survey of the general space of machine unlearning, see Nguyen et al. (2022). For a more focused survey of unlearning within the space of LLMs, see Liu et al. (2024).
@@ 7.1 Preliminaries
$\text{Pr}(\mathcal{A}(D\backslash D_{f})\in\mathcal{T})=\text{Pr}(\mathcal{U}(D,D_{f},\mathcal{A}(D))\in\mathcal{T})$
@@
$e^{-\epsilon}\leq\frac{\text{Pr}(\mathcal{U}(D,d,\mathcal{A}(D))\in\mathcal{T})}{\text{Pr}(\mathcal{A}(D\backslash d)\in\mathcal{T})}\leq e^{\epsilon}$
@@
$\text{Pr}(\mathcal{U}(D,d,\mathcal{A}(D))\in\mathcal{T})\leq e^{\epsilon}\text{Pr}(\mathcal{A}(D\backslash d)\in\mathcal{T})+\delta$
@@
$\text{Pr}(\mathcal{A}(D\backslash d)\in\mathcal{T})\leq e^{\epsilon}\text{Pr}(\mathcal{U}(D,d,\mathcal{A}(D))\in\mathcal{T})+\delta$
@@ 7.3 Unlearning in LLMs
Work on applying machine unlearning to LMs is limited but rapidly growing. Liu et al. (2024) provides an extensive survey of the current methodologies used in unlearning and how they translate to the LLM space. They also discuss related aspects such as relevant tasks and benchmarks, evaluation methods, utility preservation, and model efficiency along with discussion of LLM unlearning’s relationship to copyright and privacy protection relevant to Sections 5 and 6.
To set the stage, we can turn to Patil et al. (2023) which studied a range of extraction attacks and defenses on an LLM that has had sensitive information deleted. They focused their attack and defense framework on unlearning approaches that directly edit model weights since it is a useful unlearning approach that also factors in the possibility of whitebox attacks. In this framework, an extraction attack for a deleted data point, $A$, is successful if $A$ is contained in a candidate set $C$ of size $B$ known as the budget. Patil et al. (2023) ran their experiment on the GPT-J, Llama-2, and GPT2-XL models trained on the CounterFact Meng et al. (2022) and zsRE Levy et al. (2017) datasets. To perform the model weight editing, Patil et al. (2023) used the ROME Meng et al. (2022) and MEMIT Meng et al. (2023) techniques. Two whitebox attacks were designed which leverage the logit lens technique nostalgebraist (2020); Geva et al. (2021) that converts the hidden state from an intermediate layer into a probability distribution over the vocabulary set. The Head Projection Attack constructs the candidate set $C$ by taking the $k$ tokens at each layer with the highest probabilities. The Probability Delta Attack takes advantage of the observation that deleted data points may have inconsistent unlearning, leading to a logit lens probability that rapidly rises or falls between layers. Thus $C$ is composed by taking the $k$ tokens at each layer that had the highest difference in probability from the previous layer. Patil et al. (2023) also used a blackbox attack, Input Paraphrasing, which prompts the LLM with paraphrased versions of the original input (the paraphrase model from Krishna et al. (2023) was used). These attacks were tested against $6$ different defense methods. There were three baseline methods, The Empty Response Defense Ouyang et al. (2022), Fact Erasure Hase et al. (2023), and Error Injection De Cao et al. (2021). Three additional defenses were designed, specifically targeted towards the above designed attacks. The Head Projection Defense and Max-Entropy Defense change the weight editing objective by either subtracting the probability of the top $k$ output or maximizing entropy at each layer. The Input Rephrasing Defense addresses the rephrasing attack by adding rephrased versions of the requested deleted point to the unlearning objective. Across this framework, Patil et al. (2023) answered multiple questions. They find that the model weight editing techniques do work in providing some unlearning capabilities. However, the designed attacks were still able to extract the deleted information. In some cases, the attacks were extremely successful such as the Head Projection Attack having $89\%$ success with budget $B=20$ against GPT-J trained on zsRE edited with MEMIT. The defenses had varying degrees of success, with Head Projection and Max-Entropy providing reductions in extraction success by up to $20-40\%$. The Input Rephrasing Defense did not prove very effective. Likewise, defenses were not universally successfule, such as the Head Projection defense being ineffective against the Probability Delta attack due to it’s tailoring specifically to the Head Projection Attack.
Patil et al. (2023) concludes that while the problem of unlearning is tractable, it is still difficult to achieve with high protection from all extraction attacks. More methods will need to continue to be developed for effective unlearning to happen in the LLM space. In that spirit, here we cover different works in the LLM unlearning space sorted by the different class of methodologies used to achieve the unlearning.
@@
Efficient LOO Retraining: Kumar et al. (2022) uses an efficient LOO retraining approach by expanding upon the SISA framework by Bourtoule et al. (2020). While the SISA framework showed computational improvements over naive deletion and retraining, it is impractical for application to LLMs since the number of model parameters and data points makes it extremely computationally expensive. Kumar et al. (2022) proposed two variants of SISA that improve it for LLMs. The first is called SISA-FC. Since saving a checkpoint for every slice during the constituent model training is impossible for large text corpora, this framework starts with an LM trained on a generic corpus. The SISA framework is then only used for the task-specific fine-tuning process. While SISA-FC reduces the retraining time and computational storage, model performance generally performs better if the model is fine-tuned for the task the entire time. The second variant, known as SISA-A, addresses this by using the SISA framework for the entire training. To save memory, the SISA-A framework trains the model using the parameter efficient adapter method from Houlsby et al. (2019). Adapter modules are used to replace the model parameters and are significantly smaller, using $1-5\%$ the memory that storing all the model parameters does. This significantly reduces the memory cost of checkpointing during the SISA framework. Kumar et al. (2022) tested these two frameworks on three Glue tasks/corpora Wang et al. (2018): SST-2 Socher et al. (2013), QQP Wang et al. (2018), and MNLI Williams et al. (2018). They trained a BERT model using the different approached and measured the accuracy on these tasks, the time it took to retrain the model after a deletion request, and the amount of memory the final model takes up. SISA-A saw a small dip ($1-6\%$) in performance on the tasks but also saw a 100×100$\times$100 × improvement in retraining time. SISA-FC saw a $30\%$ dip in performance on the tasks compared to SISA-A but also was about 100×100$\times$100 × faster to retrain than SISA-A, making it way faster than the baseline approach. Exact memory numbers were not provided but Kumar et al. (2022) stated that both frameworks saw signficant reductions in memory usage. Since the SISA framework eliminates any influence the requested deleted point may have had, this approach satisfies the exact unlearning definition.
@@
Gradient-Based Updates: Jang et al. (2023) studied a gradient based solution for machine unlearning in LMs, named knowledge unlearning. Specifically, given $h_{\theta}$ and a given sequence of tokens that we would like to unlearn, $\textbf{d}=(d_{1},...,d_{T})$, $\mathcal{U}$ solves a gradient ascent problem maximizing the loss function:
@@
While leveraging In-Context Learning for unlearning purposes is effective from a privacy perspective, Muresanu et al. (2024) highlights that there can be an issue with computation cost. Specifically, ICL requires clustering the dataset to later draw training examples for the pre-pended prompt at inference time. Since deleting a given datapoint would influence the clusters, the cost of unlearning for an ICL based method is dependent on the size of the dataset. Muresanu et al. (2024) introduces their algorithm, ERASE, which uses quantized k-means Ginart et al. (2019) for their clustering. Quantized k-means introduces stability to the clusters such that with provably high probability, clusters will remain the same if a data point is deleted. Muresanu et al. (2024) showed that ERASE had similar unlearning performance to other ICUL methods and SISA unlearning while not being dependent on dataset size.
@@
7.3.1 Additional Remarks
Evaluating Unlearning Success: While these unlearning algorithms continue to be developed for LLMs, the continued question remains of how strong the privacy protection is. Hayes et al. (2024) considers the success of an approximate unlearning algorithm as whether an inference model (called a U-MIA) can distinguish between the unlearned model and a completely retrained one. More notably, Hayes et al. (2024) distinguishes between a “population U-MIA”, where one attack model is trained and used for all data points, and a “per-example U-MIA” where a dedicated attacker is instantiated for each data point. The baseline methods used to validate most unlearning methods are population U-MIAs. However, Hayes et al. (2024) adapted the LiRA attack by Carlini et al. (2022) towards this task (called U-LiRA) and found that it was highly successful in distinguishing the unlearned and retrained models. They tested this on the PaLM 2 model Anil et al. (2023), fine tuned on LIMA Zhou et al. (2023a), and unlearned using NegGrad and GradDesc Jang et al. (2023). Using this per-example U-MIA, Hayes et al. (2024) was able to identify the unlearned model over $80\%$ of the time. Given this, they argue that more stringent methods should be used to evaluate approximate unlearning methods to avoid a false sense of security.
Unlearning as a Privacy Risk: Zanella-Béguelin et al. (2020) found that model updates (such as unlearning) can possibly pose a privacy risk if an adversary has access to the versions of the model pre and post update. Through a differential analysis of the model output, a powerful membership inference attack can be performed on the model. For more details on this, see Section 4. Their attack highlights how machine unlearning is not an orthogonal concept to differential privacy, and can in some cases increase risk from a data privacy perspective.